OpenStudy (anonymous):

Solve the equation for y. Then graph the equation. 5x – 4y = 12

OpenStudy (anonymous):

You cant solve for y,Without any other equation in x and y.

OpenStudy (anonymous):

\[-4y=-5x+12\] \[y=(-5x+12)\div-7\]

OpenStudy (anonymous):

Yeah You can plot the graph for y and x as in it would be a straight line. y = (5/4)x -3. So, you have the slope = (5/4) = tan \[\theta\]

OpenStudy (anonymous):

The -7 should be -4. That gives you a value for y. Then you can use y=mx+b to plot the rest. m=-5/-4 b= 12/-4

OpenStudy (anonymous):

Yep , Agreed with you Demonfeed